OpenStudy (anonymous):

a constant F = 7.4 N net force acts for t =3.6s on a m=7.3kg object. what is the objects change in velocity

OpenStudy (anonymous):

F = M *A

OpenStudy (xishem):

Yep. You'll need F=ma and 1D kinematics equations.

OpenStudy (anonymous):

7.4 = 7.3 * a a = 7.4/7.3 = 1.01

OpenStudy (anonymous):

a = (V-u)/t

OpenStudy (anonymous):

nw find V-u